Find a matrix P such that P-1AP is a diagonal matrix. Let A =
Answer #1

\large \boldsymbol{P=\begin{bmatrix}4&-1\\1&1 \end{bmatrix}}

6 0 2 =[Y บัน + So elgon vector = for 2 5 A-XI C 너의 Veduced form.is Basson Now solve 0 01-17 Vit20 UU2 then V2 So 9gon = velt